OpenStudy (naga447):

help! What is 3 to the power of 3 over 2 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27

OpenStudy (anonymous):

\[3^{\frac{ 3 }{ 2 }}\] ?

OpenStudy (anonymous):

rules of exponents tells us we can change this to \[(3^{3})^{\frac{ 1 }{ 2 }}\] can you take over from here?

OpenStudy (naga447):

hold on let me try this.... (3^3)^1/2 3^3 = 27 27^1/2?

OpenStudy (naga447):

@binarymimic I'm not sure what happened but I don't think I did it right....

OpenStudy (anonymous):

you did it right

OpenStudy (anonymous):

27^(1/2) is the same as square root of 27

OpenStudy (naga447):

oh okay! thanks a lot!